# Resource-Constrained Federated Continual Learning: What Does Matter?

### v1

@@ III-A Experiment Setup
(a) Performance comparison of various methods in two incremental scenarios w.r.t. sufficient resources.
@@
FedAvg [1] 40.94±1.32 54.33±1.03 25.14±0.87 35.97±1.12 32.87±0.45 48.55±0.34 80.13±0.37 90.59±0.31 45.82±0.58 47.65±0.47 50.55±0.63 55.36±0.51
FedProx [12] 40.10±0.92 53.52±0.73 25.23±0.71 35.83±0.83 29.92±0.67 45.12±0.53 81.35±0.44 91.73±0.39 46.14±0.61 49.90±0.54 51.66±0.59 56.29±0.44
FedAvg+ER [42] 42.33±1.54 55.15±1.29 27.07±0.84 37.77±1.06 33.31±0.52 48.56±0.49 80.48±0.30 90.86±0.27 46.72±0.53 49.50±0.42 52.62±0.57 56.16±0.51
FedProx+ER [42] 41.07±1.07 54.53±0.98 26.97±0.78 36.10±0.96 34.08±0.79 45.54±0.66 82.06±0.49 92.68±0.51 46.21±0.73 51.66±0.67 53.58±0.68 57.18±0.56
FedCIL [14] 45.35±1.76 56.54±1.35 24.88±1.44 34.70±1.08 28.96±0.90 44.54±0.71 85.09±0.91 92.83±0.68 48.78±0.84 50.24±0.72 54.69±1.03 58.15±1.17
TARGET [13] 43.78±1.67 56.34±1.38 24.01±1.33 33.38±0.87 29.14±0.70 45.02±0.54 84.73±1.06 93.11±0.88 48.29±1.17 50.32±0.85 53.19±0.82 55.99±0.73
AF-FCL [16] 44.95±1.39 57.09±1.19 24.62±0.93 34.60±0.88 27.15±1.12 43.58±0.74 85.99±0.81 93.49±0.54 49.07±0.74 50.87±0.61 54.12±1.06 56.94±0.77
Re-Fed [15] 43.44±0.51 57.52±0.45 27.56±0.31 37.82±0.11 35.99±0.41 52.19±0.27 84.11±0.36 94.33±0.29 49.53±0.36 52.32±0.30 55.08±0.49 58.82±0.24
GLFC [17] 43.08±0.76 55.02±0.68 26.69±0.23 36.11±0.18 34.73±0.56 47.37±0.31 79.15±0.43 90.46±0.19 45.83±0.23 47.31±0.20 52.37±0.34 54.86±0.15
FOT [18] 43.74±1.02 58.24±0.78 28.43±1.14 38.57±0.95 34.23±0.74 49.52±0.61 82.97±0.67 90.40±0.55 49.35±0.98 52.09±0.81 52.18±0.53 55.11±0.60
CFeD [19] 46.54±1.13 58.23±1.78 28.37±0.46 35.43±0.86 33.05±0.30 47.98±0.62 80.23±0.11 89.13±0.22 44.26±1.69 46.78±1.64 50.13±0.53 54.89±0.92
FedWeIT [24] 41.52±1.11 54.65±0.95 26.02±0.94 36.38±0.72 34.13±0.51 49.24±0.46 81.65±0.58 91.37±0.30 46.72±0.63 48.83±0.44 51.97±0.27 56.40±0.19
(b) Performance comparison of various methods in two incremental scenarios w.r.t. extremely limited resources.
@@
FedAvg [1] 10.11±1.11 21.76±0.78 4.88±0.79 12.37±1.12 16.11±0.65 30.99±0.41 72.00±0.45 80.57±0.27 8.42±0.73 8.97±0.51 9.65±0.35 10.13±0.12
FedProx [12] 10.04±1.30 21.62±1.23 4.59±1.70 11.44±1.52 16.92±0.83 31.24±0.57 71.68±1.14 80.85±0.85 8.98±0.74 9.45±0.39 9.370.77 9.84±0.42
FedCIL [14] 7.08±1.84 17.46±1.51 3.83±1.46 9.26±1.19 4.13±1.69 7.35±1.38 54.86±1.59 62.71±1.17 5.54±0.98 7.06±0.85 7.13±1.04 7.70±0.64
FOT [18] 8.76±0.93 19.57±0.81 4.85±0.59 11.21±0.67 10.44±0.52 19.98±0.40 70.96±0.75 78.13±0.62 8.17±0.71 10.02±0.76 9.88±0.48 11.24±0.53
FedWeIT [24] 9.37±1.05 20.52±0.67 5.95±0.95 12.32±0.87 17.32±0.53 31.60±0.32 71.63±0.42 81.85±0.36 9.02±0.65 9.13±0.43 8.42±0.22 9.32±0.16
III-A2 Baselines.
In this paper, we follow the same protocols proposed by [1, 43] to set up FIbasL tasks. We evaluate our resource-constrained settings with the following baselines.
FedAvg [1]: It is a representative federated learning model which aggregates client parameters in each communication. It is a simple yet effective model for federated learning.
FedProx [12]: It is also a representative federated learning model, which is better at tackling heterogeneity in federated networks than FedAvg.
FedAvg/FedProx+ER [42]: Experience Replay, also known as Replay Buffer or Episodic Memory, is a technique primarily utilized in reinforcement learning, especially in the context of deep reinforcement learning algorithms. Here we combine the ER algorithm with the representative FL algorithms: FedAvg and FedProx.
GLFC [17]: This approach addresses the federated class-incremental learning and trains a global model by computing additional class-imbalance losses. A proxy server is introduced to reconstruct samples to help clients select the best old models for loss computation.
FedCIL [14]: This approach employs the ACGAN backbone to generate synthetic samples to consolidate the global model and align sample features in the output layer. The authors conducted experiments in the FCIL scenario, and we adopted it in our FDIL setting.
TARGET [13]: This method uses a generator to produce synthetic data to simulate the global data distribution on each client without extra data from previous tasks. The data heterogeneity is considered with catastrophic forgetting, and the authors focus on the FCIL scenario.
AF-FCL [16]: AF-FCL focuses on the data noise among previous tasks, which may harm the model performance. The NF model is proposed to quantify the credibility of previous knowledge and selectively transfer knowledge.
Re-Fed [15]: This approach proposes synergistic replay for each client to selectively cache samples based on the understanding of both local and global data distributions. Each client trains an additional personalized model to discern the importance of the score to each client.
FOT [18]: This method uses orthogonal projection technology to project different parameters into different spaces to isolate task parameters. In addition, this method proposes a secure parameter aggregation method based on projection. In the model inference stage, the method requires assuming that the boundaries of the task are known. We modified it to an automated inference of the model.
CFeD [44]: It employs public datasets as proxy data for knowledge distillation at both client and server sides. Upon introducing new tasks, clients use proxy data to transfer knowledge from previous to current models thus mitigating inter-task forgetting.
FedWeIT [24]: This method divides parameters into task-specific and shared parameters. A multi-head model based on known task boundaries is used in the original method. To ensure a fair comparison, we modify it to automate the inference of the model.
III-A3 Resource Constraints.
We analyze three common constrained resources at edge devices: Memory Buffer, Computational Budget, and Label Rate. Here, we abbreviate the memory buffer size as $M$, which means the number of the cached samples in the edge storage, and the computational budget amount as $B$ that denotes the gradient step during the model update; Label Rate as $R$, which represents the ratio of labeled data and unlabeled data. We have adopted different restriction methods for memory buffers and computational budgets for different baselines. The memory buffer is used to cache samples from previous tasks (Re-Fed [15], GLFC [17]), synthetic data (FedCIL [14], TARGET [13], AF-FCL [16]), distillation samples (CFeD [19]). The computational budget is shared by the generative model (FedCIL [14], AF-FC L[16]), personalized model (Re-Fed [15]), knowledge distillation (FOT [18], CFeD [19]). The label rate is equally employed to all baselines, which controls the ratio of the labeled data in the original data. Although we try to eliminate the gap between baseline methods other than the core algorithm design in the experiment, there are still some methods that directly depend on the backbone model itself. In our experiments, FedCIL [14] employs ACGAN for generative replay, and AF-FCL [16] needs to use a Normalized Flow model to generate the feature space for an accurate forgetting. In addition, we try to adopt the ResNet18 to the FOT [18] but achieve an unstable result. According to the available code in the original manuscript, we use AlexNet as the backbone model for the FOT [18]. The feasibility of using the ResNet network and achieving excellent performance through simple parameter tuning in the FOT algorithm is still under investigation.
III-A4 Training Details.
Unless otherwise mentioned, we employ ResNet18 [45] as the basic backbone model for most methods and use the Dirichlet distribution Dir($\alpha$) to distribute local samples to yield data heterogeneity for all tasks where a smaller $\alpha$ indicates higher data heterogeneity. Here we report the final accuracy $A(f)$ when the model finishes the training of the last streaming task and the average accuracy $\bar{A}$ across all tasks in the task streaming. Each experiment setting is run twice, and we take each run’s final 10 rounds’ accuracy and calculate the average value and standard variance. We use Adam as an optimizer with a linear learning rate schedule. We set the remaining parameters according to the values in the original open-source code, such as the weight of multiple distillation losses. All experiments are run on 8 RTX 4090 GPUs and 8*2 RTX 3090 GPUs.

### v2

@@ 3.1.4 Configurations
Table 2: Performance comparison of various methods in two incremental scenarios w.r.t. sufficient resources.
@@
FedAvg mcmahan2017communication 40.94±1.32 54.33±1.03 25.14±0.87 35.97±1.12 32.87±0.45 48.55±0.34 80.13±0.37 90.59±0.31 45.82±0.58 47.65±0.47 50.55±0.63 55.36±0.51
FedProx li2020federated 40.10±0.92 53.52±0.73 25.23±0.71 35.83±0.83 29.92±0.67 45.12±0.53 81.35±0.44 91.73±0.39 46.14±0.61 49.90±0.54 51.66±0.59 56.29±0.44
FedAvg+ER rolnick2019experience 42.33±1.54 55.15±1.29 27.07±0.84 37.77±1.06 33.31±0.52 48.56±0.49 80.48±0.30 90.86±0.27 46.72±0.53 49.50±0.42 52.62±0.57 56.16±0.51
FedProx+ER rolnick2019experience 41.07±1.07 54.53±0.98 26.97±0.78 36.10±0.96 34.08±0.79 45.54±0.66 82.06±0.49 92.68±0.51 46.21±0.73 51.66±0.67 53.58±0.68 57.18±0.56
FedCIL qi2023better 45.35±1.76 56.54±1.35 24.88±1.44 34.70±1.08 28.96±0.90 44.54±0.71 85.09±0.91 92.83±0.68 48.78±0.84 50.24±0.72 54.69±1.03 58.15±1.17
TARGET zhang2023target 43.78±1.67 56.34±1.38 24.01±1.33 33.38±0.87 29.14±0.70 45.02±0.54 84.73±1.06 93.11±0.88 48.29±1.17 50.32±0.85 53.19±0.82 55.99±0.73
AF-FCL wuerkaixi2023accurate 44.95±1.39 57.09±1.19 24.62±0.93 34.60±0.88 27.15±1.12 43.58±0.74 85.99±0.81 93.49±0.54 49.07±0.74 50.87±0.61 54.12±1.06 56.94±0.77
Re-Fed li2024efficient 43.44±0.51 57.52±0.45 27.56±0.31 37.82±0.11 35.99±0.41 52.19±0.27 84.11±0.36 94.33±0.29 49.53±0.36 52.32±0.30 55.08±0.49 58.82±0.24
GLFC dong2022federated 43.08±0.76 55.02±0.68 26.69±0.23 36.11±0.18 34.73±0.56 47.37±0.31 79.15±0.43 90.46±0.19 45.83±0.23 47.31±0.20 52.37±0.34 54.86±0.15
FOT bakman2023federated 43.74±1.02 58.24±0.78 28.43±1.14 38.57±0.95 34.23±0.74 49.52±0.61 82.97±0.67 90.40±0.55 49.35±0.98 52.09±0.81 52.18±0.53 55.11±0.60
CFeD ma2022continual 46.54±1.13 58.23±1.78 28.37±0.46 35.43±0.86 33.05±0.30 47.98±0.62 80.23±0.11 89.13±0.22 44.26±1.69 46.78±1.64 50.13±0.53 54.89±0.92
FedWeIT yoon2021federated 41.52±1.11 54.65±0.95 26.02±0.94 36.38±0.72 34.13±0.51 49.24±0.46 81.65±0.58 91.37±0.30 46.72±0.63 48.83±0.44 51.97±0.27 56.40±0.19
Table 3: Performance comparison of various methods in two incremental scenarios w.r.t. extremely limited resources.
@@
FedAvg mcmahan2017communication 10.11±1.11 21.76±0.78 4.88±0.79 12.37±1.12 16.11±0.65 30.99±0.41 72.00±0.45 80.57±0.27 8.42±0.73 8.97±0.51 9.65±0.35 10.13±0.12
FedProx li2020federated 10.04±1.30 21.62±1.23 4.59±1.70 11.44±1.52 16.92±0.83 31.24±0.57 71.68±1.14 80.85±0.85 8.98±0.74 9.45±0.39 9.370.77 9.84±0.42
FedCIL qi2023better 7.08±1.84 17.46±1.51 3.83±1.46 9.26±1.19 4.13±1.69 7.35±1.38 54.86±1.59 62.71±1.17 5.54±0.98 7.06±0.85 7.13±1.04 7.70±0.64
FOT bakman2023federated 8.76±0.93 19.57±0.81 4.85±0.59 11.21±0.67 10.44±0.52 19.98±0.40 70.96±0.75 78.13±0.62 8.17±0.71 10.02±0.76 9.88±0.48 11.24±0.53
FedWeIT yoon2021federated 9.37±1.05 20.52±0.67 5.95±0.95 12.32±0.87 17.32±0.53 31.60±0.32 71.63±0.42 81.85±0.36 9.02±0.65 9.13±0.43 8.42±0.22 9.32±0.16
3.2 Overview Performance under Resource Constraints
In this section, we systematically investigate the influence of resource constraints on existing FCL methods and typical techniques studied in the FCL literature.
With Sufficient Resources. In Table 2, we test typical FCL methods under sufficient resource conditions. Notably, both FedCIL and AF-FCL methods excel, surpassing most established baselines on CIFAR-10 and Digit-10. This superior performance underscores the effectiveness of their generative models, which can effectively train on relatively simple datasets and generate high-quality samples for replay. However, as we delve into more complex and larger-scale datasets like CIFAR-100 and Tiny-ImageNet, we observe a significant drop in performance for both FedCIL and AF-FCL. This decline highlights the limitations of their generative models when faced with greater data diversity and complexity. In the context of federated domain-incremental learning, where the data distribution changes over time, GLFC demonstrates an interesting behavior. When there are no new incremental sample classes, it reverts to the FedAvg algorithm. On the other hand, Re-Fed consistently achieves reasonable performance across all datasets and scenarios. Its synergistic replay strategy against data heterogeneity seems to provide a robust foundation for FCL.
With Limited Resources. The experiments in Table 3 conducted under extremely limited resource conditions starkly contrast to the results obtained with sufficient resources. By simultaneously restricting three key types of training resources, we observe a significant decline in the performance of all methods. This underscores the critical importance of these resources for the effectiveness of federated continual learning techniques. To better understand the impact of these resource constraints, we conducted experiments from two perspectives. First, we analyzed how each resource limitation affects the performance of the various methods. This allowed us to identify the most sensitive components and bottlenecks that hinder the effectiveness of the FCL techniques under limited conditions. Second, we explored the interactions between different resource constraints and how they compound the challenges faced by the FCL methods. By understanding these interactions, we can gain insights into designing more resource-efficient FCL techniques. A detailed analysis will be provided in the next section.
3.2.1 Ablation Study on Three Training Resources
Figure 1: Ablation study on three training resources.
1. Does the Memory Buffer matter? As shown in Fig.1(a), we select six methods that need extra memory buffer to cache previous samples/synthetic samples/auxiliary datasets. Specifically, when allocated a modest amount of memory buffer, these methods experience a marked improvement in their performance. This enhancement becomes particularly pronounced for methods that leverage generative models and distilled data for datasets with simple features. However, as the complexity of the datasets increases, posing challenges such as a wider diversity of classes and domains, the performance of these methods, particularly those reliant on generative modeling, undergoes a noticeable decline. This underscores the limitations of solely relying on synthetic data generation or distillation in complex learning environments. In contrast, methods that employ straightforward sample caching exhibit a more resilient performance profile. They maintain a relatively stable performance, demonstrating their robustness against dataset complexity. Nevertheless, as the number of cached samples surpasses a certain threshold, the model’s performance gradually approaches an upper limit, indicating diminishing returns from further sample accumulation. This plateauing effect underscores the need for a delicate balance in managing the memory buffer, aiming to maximize performance gains without incurring unnecessary computational cost.
2. Does the Computational Budget matter? As depicted in Fig.1(b), we carefully allocated varying computational budgets for each dataset, taking into account the maximum allowable limit specified in Table 5. It is evident from the results that when the computational budgets assigned to the various methods are restricted, the overall performance of the models inevitably suffers. This underlines the critical role that computational resources play in ensuring optimal model performance. However, certain methods demonstrate remarkable resilience, notably FedAvg and FOT. Even with limited computational budgets, they maintain satisfactory performance levels, highlighting their efficiency and effectiveness in resource-constrained environments. On the other hand, when the computational budget is abundant, allowing for unrestricted resource allocation, these other methods display the potential for higher performance ceilings. Given sufficient resources, they may achieve performance that surpasses that of FedAvg and FOT. Nevertheless, the practical implications of such high resource requirements must be carefully considered, as they may not always be feasible or cost-effective in real-world applications.
3. Does the Label Rate matter? In Fig.1(c), we analyze the relationship between the label rate and the model performance. Our analysis reveals a compelling trend: when the label rate falls below 70%, a strategic increase in the proportion of labeled data within the dataset is a potent catalyst for enhancing the performance of most model approaches. However, the Digit-10 dataset stands out as a unique case. Despite boasting a vast quantity of data, the simplicity of its feature information, characterized by single-channel data, renders it less reliant on an extensive labeled dataset for optimal performance. Consequently, even with a relatively modest amount of labeled data, the model can achieve promising results, underscoring the importance of considering the dataset’s inherent characteristics when evaluating the label rate’s impact. This observation underscores the nuanced interplay between the label rate, dataset complexity, and model performance. While a higher label rate generally leads to improved performance, the extent of this improvement can vary significantly depending on the dataset’s characteristics and the model being employed.
Conclusion. Existing methods have encountered limitations in achieving optimal results within resource-constrained environments. Notably, deploying a modestly sized memory buffer has emerged as a crucial factor in significantly augmenting model performance. This enhancement underscores the importance of an adequately provisioned yet economically efficient memory allocation strategy, as an overly generous buffer does not necessarily translate into commensurate gains in performance. A similar trend is observed with the label rate, emphasizing the need for a balanced approach in labeling data to maximize model efficacy. Moreover, the influence of the computational budget on FCL performance is intimately tied to the algorithmic design itself. Simple yet effective methods such as FedAvg demonstrate remarkable resilience, maintaining robust performance even with stringent computational constraints. This finding underscores the potential of streamlined FCL algorithms in addressing real-world challenges posed by limited resources. Experimental evidence highlights the imperative of selecting FCL methods tailored to the specific training resources. Researchers can achieve more effective learning outcomes by aligning algorithmic complexity with computational and memory resources.
Figure 2: Analysis of three typical FCL techniques.
3.2.2 Analysis of Typical FCL Techniques
1. Does the Sample Caching matter? To gain a deeper understanding of the efficacy of sample caching technology within limited resources, we impose restrictions on the memory buffer size and the label rate illustrated in Fig.2(a). Regarding the sample caching technique, the label rate is pivotal in determining the quantity and quality of samples available for local training. Specifically, a higher label rate translates into more high-quality, labeled samples within the dataset, significantly enhancing the learning process. The Re-Fed algorithm, in particular, introduces an innovative approach that leverages collaborative storage to intelligently select and retain the most critical samples when both the label rate and memory buffer are limited. This strategy is particularly effective in boosting model performance by focusing on the most informative data points, even under stringent resource constraints. Furthermore, as the label rate and memory buffer size vary, they jointly influence the performance of the sample caching technique in complex ways. The label rate not only dictates the overall quality of the data but also serves as an upper bound on the number of samples that can be feasibly stored within the given memory buffer. Consequently, the effectiveness of the sample caching is intimately tied to the interplay between these two training resources. Optimizing the allocation and utilization of these resources is thus crucial for maximizing the benefits of sample caching with limited resources.
3. Does the Knowledge Distillation matter? Knowledge distillation is a prevalent technique in mitigating catastrophic forgetting within existing FCL methods. While this approach effectively preserves key information across learning tasks, it necessitates the utilization of additional distillation datasets, whether publicly accessible or synthetically generated. Unfortunately, this requirement introduces additional complexity layers, including managing both computational overhead and memory buffer capacity. We primarily focus on limiting the computational budget, as it is often the primary factor constraining the scalability and speed of the distillation process. The Fig.2(c) illustrates that the computational budget is the primary limiting factor affecting the performance of the corresponding knowledge distillation methods. On the other hand, the memory buffer, though still a consideration, has a relatively minor impact on the overall performance of the knowledge distillation technique. This observation highlights the efficiency of the distillation process, where even a modest quantity of distilled data samples can achieve satisfactory performance gains. By efficiently managing the memory buffer to accommodate these necessary samples, we can maximize the benefits of knowledge distillation without incurring excessive memory overhead. In summary, our approach to knowledge distillation in FCL involves a careful balancing act between computational budget and memory buffer, with a greater emphasis on optimizing the former to ensure the technique’s practicality and performance within the confines of federated and continual learning environments.
Discussion about the Network Extension. In this section, we do not perform additional experimental investigations specifically tailored for the Network Extension method. The main reason is that prevalent network extension methodologies typically presuppose knowledge of task boundaries, implying that the task ID for incoming data is accessible during inference. When this impractical assumption is relaxed, and we analogize to other methodologies that infer the category of the sample without prior task information, these extensions fail to uphold their erstwhile promising performance, as evidenced in Tables 2 and 3.
Conclusion. Three FCL strategies differ in strengths and limitations under varying resource conditions. With ample computational power and memory, sample caching preserves representative previous samples to combat forgetting, while data synthesis generates pseudo samples without storing real data – albeit requiring significant computational resources for effective rehearsal. Conversely, knowledge distillation demonstrates resource efficiency by maintaining performance with minimal distilled samples, bypassing large memory buffers. Resource availability thus dictates strategy selection: sample caching and data synthesis excel in resource-rich environments, while knowledge distillation becomes optimal under constraints. Future research should prioritize reducing synthesis computation, optimizing cached memory usage, and enhancing distillation efficiency to address growing demands for resource-constrained FCL implementations.
4 Evaluation & Future Research Direction
Table 4: Performance comparison of various methods with a different order of domain tasks. We report the reverse order compared to the Table 2.
Dataset Metric FedAvg FedCIL Re-Fed FOT
Digit10 $A(f)$ 75.84±0.36 79.96±1.12 79.63±0.53 74.45±0.82
$\bar{A}$ 83.62±0.41 87.98±0.88 88.23±0.49 83.91±0.76
Office31 $A(f)$ 36.65±0.52 41.90±1.36 42.71±0.44 40.29±0.94
$\bar{A}$ 41.47±0.49 44.18±0.95 45.05±0.39 43.86±0.73
In this section, we delve deeper into the analysis of existing FCL methodologies by conducting extensive quantitative experiments that focus on two crucial aspects: the number of tasks within Class-IL datasets in Figure. 3 and the order of tasks across various Domain-IL datasets in Table 3. By systematically varying these parameters, we aim to gain a nuanced understanding of how these factors impact the performance of FCL methods. Firstly, we observe that as the number of tasks in the Class-IL datasets escalates, the performance of all the evaluated methods noticeably declines. Nevertheless, despite this performance decrement, the general trends exhibited by the different methods remain comparable, suggesting that the fundamental challenges they face are similar. Furthermore, our experiments reveal that the order in which domain tasks are presented to the model can significantly influence its ultimate performance within Domain-IL scenarios. This finding emphasizes the importance of task scheduling and curriculum design in FCL systems, as certain sequences may be more conducive to learning than others. However, even with this variability in performance due to task order, each method maintains its distinctive strengths and tendencies, demonstrating its resilience and adaptability to different contexts.
Moreover, although there is a lot of research on FCL, these studies mainly focus on transferring CL algorithms in centralized settings to alleviate catastrophic forgetting, ignoring the issue of training resource overhead. Based on this, we shed light on the future research direction in Appendix C.
5 Conclusion
This paper has presented a tutorial on FCL under resource constraints. Firstly, we begin with an introduction to the existing FCL methods and three training resources on distributed devices that influence the performance of the methods. Then, we conduct experiments using ablation studies on each training resource and analyze the effect of four typical FCL techniques on alleviating catastrophic forgetting in FCL. Finally, we discuss future research directions in resource-constrained FCL.
Acknowledgments and Disclosure of Funding
This work is supported by the National Key Research and Development Program of China under grant 2024YFC3307900; the National Natural Science Foundation of China under grants 62376103, 62302184, 62436003 and 62206102; Major Science and Technology Project of Hubei Province under grant 2024BAA008; Hubei Science and Technology Talent Service Project under grant 2024DJC078; Ant Group through CCF-Ant Research Fund; and Fundamental Research Funds for the Central Universities under grant YCJJ20252319. The computation is completed in the HPC Platform of Huazhong University of Science and Technology.
A Background and Related Work
Federated Learning. Federated Learning (FL) is an approach to developing a unified global model by combining models trained on local, private datasets from multiple clients liao2024swiss ; liao2025fedbrb ; wang2025better ; meng2024improving . A notable FL framework, FedAvg mcmahan2017communication , enhances the global model by averaging the parameters of these locally trained models. Nevertheless, conventional FL methods, such as FedAvg, encounter difficulties when dealing with data heterogeneity, where client datasets are not independently and identically distributed (Non-IID), leading to reduced model performance huang2022learn ; huangself ; huangfedbg ; qi2023cross . To address the Non-IID problem in FL, an optimization technique incorporating a proximal term was introduced in li2020federated to alleviate the impact of diverse and Non-IID data distributions among devices. Another strategy, federated distillation zhang2021parameterized ; zhu2021data , focuses on transferring knowledge from multiple local models to the global model by aggregating solely the soft predictions produced by each model. The authors of lin2020ensemble presented a knowledge distillation technique that employs unlabeled training data as a surrogate dataset. However, these methods are tailored to address static data with spatial heterogeneity and do not account for the challenges presented by temporally heterogeneous streaming tasks in FL.
Continual Learning Continual Learning (CL) is a machine learning paradigm that enables models to learn sequentially from a stream of tasks while preserving knowledge acquired from earlier tasks van2019three ; zou2024compositional ; zhou2024delve . CL methods can be broadly categorized into three primary approaches: replay-based rebuffi2017icarl ; fedus2020revisiting , regularization-based Jung2020ContinualLW ; kemker2018measuring , and parameter isolation techniques von2019continual ; pham2021dualnet . Replay-based methods involve retaining a subset of previous task samples to maintain knowledge when learning new tasks. Regularization-based methods prevent the overwriting of existing knowledge by applying constraints to the loss function during the learning of new tasks. Parameter isolation methods usually introduce extra parameters and computational resources to facilitate the learning of new tasks. In this work, we concentrate on federated continual learning, which integrates elements of both federated and continual learning.
Federated Continual Learning Federated Continual Learning (FCL) is introduced to tackle the challenge of learning from a sequence of tasks on each client, focusing on adapting the global model to new data while preserving knowledge from previous data wang2024federated ; wang2024traceable ; yu2024overcoming . Although important, FCL has only recently started to receive attention, with yoon2021federated serving as a seminal work in this area. This study concentrates on Task-Incremental Learning (Task-IL), which necessitates distinct task IDs at inference time and employs dedicated task-specific masks to improve personalized performance. Concurrently, research efforts like yu2024personalized consider the multi-granularity representation of knowledge, fostering the integration of spatial-temporal knowledge in Federated Continual Incremental Learning (FCIL). The authors of li2024pfeddil align local domain knowledge with a dynamic network to strike a balance between resource usage and model performance. Other works, such as ma2022continual , employ a proxy dataset and leverage knowledge distillation at both the server and client levels. In contrast, dong2022federated ; dong2023no simplifies the problem by assuming clients have sufficient storage to store and share old examples, deviating from the conventional FL framework. Additionally, studies like dong2023federated ; jin2024fl investigate FCL in applications beyond image classification.
B Problem Formulation
Federated Learning. In the context of Federated Learning (FL), a common problem involves collaboratively training a global model across $K$ clients. Each client $k$ has exclusive access to its local private dataset, denoted as $\mathcal{T}_{k}=\{x^{(i)}_{k},y^{(i)}_{k}\}$, where $x^{(i)}_{k}$ represents the $i$-th data sample, and $y^{(i)}_{k}\in\{1,2,\ldots,C\}$ is the associated label for $C$ classes. The size of dataset $\mathcal{T}_{k}$ is indicated by $|\mathcal{T}_{k}|$. The global dataset is the aggregate of all local datasets, i.e., $\mathcal{T}=\{\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{K}\}$, with $\mathcal{T}=\sum_{k=1}^{K}\mathcal{T}_{k}$. The goal of the FL system is to train a global model $w$ that minimizes the overall empirical loss across the entire dataset $D$:
$\text{where}\ \mathcal{L}_{k}(w)=\frac{1}{|\mathcal{T}_{k}|}\sum_{i=1}^{|\mathcal{T}_{k}|}\mathcal{L}_{CE}(w;x_{i}^{k},y_{i}^{k}).$ (1)
Here $\mathcal{L}_{k}(w)$ signifies the local loss for client $k$, and $\mathcal{L}_{CE}$ is the cross-entropy loss function, which quantifies the discrepancy between the predicted and actual labels.
Continual Learning. In a typical Continual Learning (CL) scenario (outside of a federated context), a model is trained on a sequence of tasks $\{\mathcal{T}^{1},\mathcal{T}^{2},\ldots,\mathcal{T}^{n}\}$, where $\mathcal{T}^{t}$ represents the $t$-th task. Each task $\mathcal{T}^{t}=\sum_{i=1}^{N^{t}}(x^{(i)}_{t},y^{(i)}_{t})$ comprises $N^{t}$ pairs of data samples $x^{(i)}_{t}\in\mathcal{X}^{t}$ and their corresponding labels $y^{(i)}_{t}\in\mathcal{Y}^{t}$. The domain space and label space for the $t$-th task are denoted by $\mathcal{X}^{t}$ and $\mathcal{Y}^{t}$, respectively, with $|\mathcal{Y}^{t}|$ classes. The total class set across all tasks is $\mathcal{Y}=\bigcup_{t=1}^{n}\mathcal{Y}^{t}$, and similarly, the total domain space is $\mathcal{X}=\bigcup_{t=1}^{n}\mathcal{X}^{t}$. This paper focuses on two CL scenarios: (1) Class-Incremental Task: All tasks share the same domain space ($\mathcal{X}^{1}=\mathcal{X}^{t}$ for all $t\in[n]$), but the number of classes may vary ($\mathcal{Y}^{1}\neq\mathcal{Y}^{t}$ for all $t\in[n]$). (2) Domain-Incremental Task: All tasks have the same number of classes ($\mathcal{Y}^{1}=\mathcal{Y}^{t}$ for all $t\in[n]$), but the domain and data distribution change over tasks ($\mathcal{X}^{1}\neq\mathcal{X}^{t}$ for all $t\in[n]$).
Federated Continual Learning. We combine the CL with the federated setting. Our objective is to train a global model for $K$ clients, where each client $k$ can only access its local sequence of tasks $\{\mathcal{T}_{k}^{1},\mathcal{T}_{k}^{2},\ldots,\mathcal{T}_{k}^{n}\}$. When the $t$-th task arrives, the goal is to train a global model $w^{t}$ across all $t$ tasks, denoted by $\mathcal{T}^{t}=\{\sum_{n=1}^{t}\sum_{k=1}^{K}\mathcal{T}_{k}^{n}\}$. This can be formulated as:
Here $\mathcal{L}_{CE}$ is the cross-entropy loss function used to measure the discrepancy between predictions and true labels.